# Neural variational inference for cutting feedback during uncertainty propagation

### v1

@@ 2.2 Variational Inference
Variational inference (VI) provides a practical alternative by approximating posterior distributions via optimization (Blei et al.,, 2017). Given data $D$ with model specified by some parameters $\phi$, VI seeks a distribution $q(\phi)$ within a variational family $\mathcal{Q}$ that minimizes the KLD to a posterior distribution $p(\phi\,|\,D)$, to get the optimal solution:
${q}^{*}(\phi)=\underset{q(\phi)\in\mathcal{Q}}{\arg\min}\,\,\mathrm{KL}\,\!({q}(\phi)\parallel p(\phi\,|\,D)).$ (4)
If no restriction is placed on the family $\mathcal{Q}$ of distributions, then the solution $q^{*}(\phi)$ is exactly the true posterior distribution $p(\phi\,|\,D)$. In practice, one places some parametric assumptions on the class $\mathcal{Q}$ to obtain an optimizer that approximates the true posterior distribution.
A fundamental contribution in cut-Bayes or cutting feedback methods that enabled the use of variational inference for this problem is the result (Lemma 4.1) of Yu et al., (2023). Considering the setup of Section 2.1 with $\phi=(\theta,\eta)$, the result proved that when $\mathcal{Q}$ is restricted to all distributions $q(\theta,\eta)$ such that $q(\eta)=p(\eta\,|\,D_{1})$, the optimization in (4) yields the joint cut-posterior (2) as the solution. Formally,
@@
Thus, the joint cut-posterior can be characterized as the variational solution, minimizing KLD to the joint full Bayes posterior under the constraint of no feedback from $D_{2}$ to $\eta$. In their implementation, they used two variational approximations, decomposing $q(\theta,\eta)=q(\theta\,|\,\eta)\,q(\eta)$, using parametric families for both, i.e., $q(\theta\,|\,\eta)=q_{\lambda_{2}}(\theta\,|\,\eta)$ and $q(\eta)=q_{\lambda_{1}}(\eta)$. To estimate $q^{*}(\eta)$, the KLD to $p(\eta\,|\,D_{1})$ was minimized with respect to $\lambda_{1}$ to obtain $\hat{q}^{*}(\eta)=q_{\hat{\lambda}_{1}}(\eta)$. Then $q_{\hat{\lambda}_{1}}(\eta)$ is plugged into the second optimization (5) to estimate the minimizing parameters $\hat{\lambda}_{2}$. The final cut-posterior estimate is then given by $q_{\hat{\lambda}_{2}}(\theta\,|\,\eta)\,q_{\hat{\lambda}_{1}}(\eta)$. Similar ideas were used in Smith et al., (2025) in the context of copula models.
@@ 2.3 Normalizing Flows
The effectiveness of variational inference depends on the choice of variational family. Conventional choices such as mean-field approximations, i.e., assuming $q(\phi)=\prod_{i}q_{i}(\phi_{i})$, often fail to capture complex posterior structures. Normalizing flows provide a flexible alternative by constructing complex densities through a sequence of invertible and differentiable transformations of a simple base distribution, and have been widely explored in variational inference (Rezende and Mohamed,, 2015).
The basic idea of normalizing flows is as follows. If $F$ is the cumulative distribution function (cdf) of a scalar $\theta$, then there always exists a function $f$ such that $\theta\overset{d}{=}f(Z)$, where $Z$ is some base distribution. The function $f$ is explicitly obtained from the inverse cdf transformation, e.g., if $Z\sim N(0,1)$, $f(Z)=F^{-1}(\Phi(Z))$ where $\Phi$ is the standard Normal cdf. Hence, the task of approximating the distribution of $\theta$ becomes equivalent to approximating the function $f$. As neural networks are universal approximators of many function classes, it thus suffices to model $f$ using a suitable class of neural networks $g_{\lambda}$ and estimate the neural parameters (weights and biases) $\lambda$ using the variational (KL) loss. Normalizing flows build on this core idea, but the specific choices to model $f$ vary widely. We will discuss some common flow classes when introducing our method and studying its theoretical properties. A more comprehensive review of normalizing flows and their extensions can be found in (Kobyzev et al.,, 2020; Papamakarios et al.,, 2021).
@@ 3 Neural Variational Inference for Cut-Bayes
In this section, we propose a variational inference method for cutting feedback that $(a)$ can directly use samples of $\eta$ from the upstream posterior $p(\eta\,|\,D_{1})$ without requiring access to the raw data and model; and $(b)$ uses normalizing flows to non-parametrically model the conditional cut-posterior of $\theta$ given $\eta$.
Recall from (5) that the cut-posterior is the solution to the variational optimization with the variational family $q(\theta,\eta)$ constrained to have $q(\eta)=p(\eta\,|\,D_{1})$ to prevent feedback. As $q(\eta)$ is fixed, we make two observations. First, the optimization in (5) is essentially only over the class of conditional distributions $q(\theta\,|\,\eta)$. Also, under this constraint of $q(\eta)=p(\eta\,|\,D_{1})$, the KL loss function in (5) can be written as $\mathrm{KL}\,\!({q}(\theta,\eta)\parallel p(\theta,\eta\,|\,D_{1},D_{2}))=\mathbb{E}_{\eta\sim p(\eta\,|\,D_{1})}\big[\mathrm{KL}\,\!(q(\theta\,|\,\eta)\parallel p(\theta\,|\,\eta,D_{2}))\big]$.
@@
Thus, finding the joint cut-posterior essentially reduces to estimating the collection of conditional distributions $\{q(\theta\,|\,\eta)\}_{\eta}$. No approximations to $p(\eta\,|\,D_{1})$ is needed. In practice, we will typically have samples $\eta_{1},\ldots,\eta_{N}$ from $p(\eta\,|\,D_{1})$. As long as the samples are iid or from an ergodic MCMC, we can approximate the expectation term in (6) with the sample average, leading to the estimate
@@
Equations (6) and (7) illustrate how the joint cut-posterior can be estimated without access to the upstream data $D_{1}$ and the model used in upstream analysis, as long as the samples $\{\eta_{i}\}$ are available. Even if $D_{1}$ and its model are available, there is no need to approximate $p(\eta\,|\,D_{1})$ using the additional variational inference step, as done in Yu et al., (2023); Smith et al., (2025). One can run a standard MCMC, to estimate $p(\eta\,|\,D_{1})$ and use the posterior samples $\{\eta_{i}\}$ in (7). Once the downstream analysis model likelihood $p(D_{2}\,|\,\theta,\eta)$ and the prior $p(\theta\,|\,\eta)$ are chosen, the only unknown in (7) is the class of conditional distributions $q(\theta\,|\,\eta)$. In the next section, we propose using normalizing flows that can approximate any conditional family, leading to a provably accurate estimate of the joint cut-posterior.
@@ 3.1 Normalizing Flows for Conditional Distribution
We propose modeling $q(\theta\,|\,\eta)$ using conditional normalizing flows, which approximate complex distributions using transformations of simple base distributions. Specifically, if $\theta\in\mathbb{R}^{d}$, we consider a class of conditional distributions $q_{\vartheta}(\theta\,|\,\eta)$ that corresponds to the law $\theta\,|\,\eta\sim T_{\vartheta}(\eta,Z)$. Here $T_{\vartheta}(\eta,Z)$ is a class of neural network-based functions (choices discussed later) with parameters $\vartheta$ taking values in some space $\Theta$, and $Z\in\mathbb{R}^{d}$ is a latent variable of the same dimension as $\theta$ and with a base distribution $p(Z)$ that is easy to sample from. A common choice is the standard normal distribution, e.g., $Z\sim N(0,I_{d})$. However, if the true posterior is heavy-tailed, then Cauchy or Student’s $t$ may be a better base. See Section 4 for theoretical insights on the choice of the base.
@@
By drawing $Z_{i}\overset{\text{iid}}{\sim}p(Z)$, we can have samples $\{\eta_{i},T_{\hat{\vartheta}}(\eta_{i},Z_{i})\}$ from this cut-posterior. We call our method NeVI-Cut (neural variational inference for cut-Bayes). Note that NeVI-Cut is completely agnostic to what upstream data $D_{1}$ and model were used to obtain the samples $\{\eta_{i}\}$ as long as averages using this sample approximate population averages with respect to $p(\eta\,|\,D_{1})$. It also works with any choice of base distribution, any dimension, and support set of the parameters $\theta$ and $\eta$, and any class of normalizing flows. The next section presents some important classes of normalizing flows that we use in practice. The theoretical justification for their use is presented in Section 4.
3.2 Choices of Flows
We first consider the rational-quadratic neural spline flows (RQ-NSF) (Durkan et al.,, 2019) with two widely used flow architectures: an autoregressive architecture referred to as RQ-NSF(AR) (Papamakarios et al.,, 2017), and a coupling architecture referred to as RQ-NSF(C) (Dinh et al.,, 2016; Kingma and Dhariwal,, 2018).
In RQ-NSF(AR), each transformation $f_{k}$ of the function $T$ is parameterized by an autoregressive neural network. Let $Z=(Z_{[1]},...,Z_{[d]})\in\mathbb{R}^{d}$ denote the base variable. In this architecture, each output dimension $Z_{[i]}^{(k)}$ at layer $k$ only depending on the previous components $z_{[{<i}]}^{(k-1)}$ and the conditional input $\eta$. Specifically, the transformation at flow layer $k$ is defined as
$Z_{[i]}^{(k)}=f_{k}\left(Z_{[i]}^{(k-1)};\phi_{k}\left(Z_{[<i]}^{(k-1)},\eta\right)\right),$
where $f_{k}$ is rational quadratic spline transformation (Durkan et al.,, 2019), and $\phi_{k}$ is a masked conditioner neural network that outputs the spline parameters based on previous latent variables and $\eta$. Since each one-dimensional RQS is strictly monotone, invertibility is guaranteed, and the autoregressive composition preserves this property across all dimensions. The autoregressive structure yields a lower triangular Jacobian, making the log-determinant straightforward to compute. To ensure all coordinates are eventually updated, permutation layers (e.g., reversing or rotating the variable order) are inserted between autoregressive transformations.
RQ-NSF(C) is a coupling-based flow that applies monotonic rational quadratic splines transformation to a subset of variables while leaving others unchanged. For each transformation $f_{k}$, the input $Z^{(k-1)}$ is partitioned into two disjoint subsets $Z_{[\alpha]}^{(k-1)}$ and $Z_{[\beta]}^{(k-1)}$. The variables in $Z_{[\alpha]}^{(k-1)}$ are unchanged, the monotone RQS transformations are applied to $Z_{[\beta]}^{(k-1)}$ according to
$Z_{[\beta]}^{{(k)}}=f_{k}\left(Z_{[\beta]}^{(k-1)};\psi_{k}\left(Z_{[\alpha]}^{(k-1)},\eta\right)\right),$
where $\psi_{k}(\cdot)$ is a neural network that outputs spline parameters (e.g., bin widths, heights, and slopes) conditioned on the retained subset and conditional input, and together $Z^{(k)}=\left(Z_{[\alpha]}^{(k-1)},Z_{[\beta]}^{{(k)}}\right)$ is passed to the next layer. The partition $(\alpha,\beta)$ is specified by a binary mask, following the coupling design of RealNVP (Dinh et al.,, 2016), and alternated or permuted across layers as in Neural Spline Flows (Durkan et al.,, 2019) so that all coordinates are eventually updated. Because the variables in $\alpha$ are unchanged, the Jacobian has a block-triangular form with an identity block for $\alpha$ and a diagonal block for $\beta$, making the log-determinant efficient to compute (Dinh et al.,, 2016).
In addition to spline-based flows, we also consider Unconstrained Monotonic Neural Networks (UMNN; Wehenkel and Louppe,, 2019), which construct invertible transformations by integrating positive neural network outputs. We extend this idea to a conditional setting. Similar to the original UMNN, our construction specifies the variational distribution $q(\theta\,|\,\eta)$ induced by the monotone transformation conditioned on input $\eta$, defined as
$\theta=T_{\vartheta}(\eta,Z)=a_{\vartheta_{a}}(\eta)+\int_{0}^{Z}\exp\left(g_{\vartheta_{g}}(\eta,t)\right)\,dt.$ (11)
Here $a(\cdot)$ is a neural network with parameters $\vartheta_{a}$ that shifts the transformation, and $g(\cdot)$ is a neural network with parameters $\vartheta_{g}$ whose output is passed through an exponential function and models the log-gradient of $T_{\vartheta}$ with respect to $Z$, and $\vartheta=(\vartheta_{a},\vartheta_{g})$. This guarantees that the derivative of $T_{\vartheta}$ with respect to $Z$ is strictly positive and hence that the transformation is monotone and invertible.
@@ 3.3 Triply-Stochastic Algorithm
The computational challenges in calculating the objective in (9) can be three-pronged. First, a large number of samples $\eta_{i}$ may be available from the upstream analysis. Second, the expectation with respect to $Z$ is typically not available in closed form. Third, the likelihood $p(D_{2}\,|\,\theta,\eta)$ might be slow to compute especially for large datasets. We address these challenges by leveraging three sources of stochasticity.
Mini-batching or stochastic gradient descent is widely used to speed up neural network optimization. The main idea is that terms in the objective function involving sums or averages over many variables can be replaced by averages (or appropriately scaled sums) over smaller subsets of these variables. These subsets are referred to as mini-batches. Only one batch is used at a time, and the mini-batches are cycled throughout the optimization. Stochastic gradient descent is an extreme case with a mini-batch size of one. Each of these three challenges has a scalable solution using such stochastic approximations. At each iteration, we can replace the average over all $N$ samples of $\eta$ with a small minibatch $\mathcal{B}_{\eta}$ of size $N_{\eta}$. Similarly, the expectation with respect to $Z$ is replaced by an average of $N_{Z}$ iid observations $\{Z_{j}\}$ sampled from $p(Z)$. The doubly stochastic approximation of (9) is thus given by
$\hat{\mathcal{L}}_{DS}(\vartheta)=\frac{1}{N_{\eta}}\frac{1}{N_{Z}}\sum_{i\in\mathcal{B}_{\eta}}\sum_{j=1}^{N_{Z}}\Big[$ $\log|\det J_{z}T_{\vartheta}(\eta_{i},Z_{j})|+$ (12)
@@
Finally, if the downstream data $D_{2}$ consists of iid units $Y_{1},\ldots,Y_{n}$, then we have $p(D_{2}\,|\,\theta,\eta)=\prod_{i=1}^{n}p(Y_{i}\,|\,\theta,\eta)$. And $\log p(D_{2}\,|\,\theta,\eta)$ can be approximated as $\frac{n}{N_{D}}\sum_{k\in\mathcal{B}_{Y}}\log p(Y_{k}\,|\,\theta,\eta)$. This gives the triply stochastic approximation
$\hat{\mathcal{L}}_{TS}(\vartheta)=\frac{1}{N_{\eta}}\frac{1}{N_{Z}}\sum_{i\in\mathcal{B}_{\eta}}\sum_{j=1}^{N_{Z}}\Big[$ $\log|\det J_{z}T_{\vartheta}(\eta_{i},Z_{j})|+$ (13)
$\frac{n}{N_{D}}\sum_{k\in\mathcal{B}_{Y}}\log p\big(Y_{k}\,|\,T_{\vartheta}(\eta_{i},Z_{j}),\eta_{i}\big)+\log p\!\left(T_{\vartheta}(\eta_{i},Z_{j})\,|\,\eta_{i}\right)\Big].$
@@
Input: Upstream posterior samples $\{\eta_{i}\}_{i=1}^{N}$, downstream data $D_{2}$, functional form of downstream likelihood $p(D_{2}\,|\,\theta,\eta)$ and downstream prior $p(\theta\,|\,\eta)$, base distribution $p(Z)$, maximum number of iterations $S$, patience $S_{\text{patience}}$ (number of steps before early stopping), neural class $T$ (e.g., Rational Quadratic Splines or Unconstrained Monotonic Neural Network), learning rate $\alpha$.
@@
  Estimate Monte Carlo objective:
$\hat{\mathcal{L}}_{Z}(\vartheta)=\frac{1}{M}\sum_{i=1}^{N}\big[\log|\det J_{z}T_{\vartheta}(\eta_{i},Z_{i})|+\log p\!\left(D_{2}\mid\theta_{i},\eta_{i}\right)+\log p\!\left(\theta_{i}\mid\eta_{i}\right)\big].$
@@
Output: Trained conditional flow $\hat{q}(\theta\,|\,\eta)\overset{\text{d}}{=}T_{\hat{\vartheta}}(\eta,Z),Z\sim p(Z)$, samples from the cut-posterior $\{\eta_{i},\widetilde{\theta}_{i}\}_{i=1}^{N}$.
In practice, all three sources are not required simultaneously, and the level of stochasticity, i.e., choices of the minibatch sizes $N_{\eta},N_{Z},N_{D}$ can be adjusted depending on computational trade-offs. For example, if the data are not iid or not large, mini-batching of the likelihood is not recommended, and the third source can be omitted. In our experiments, to ensure comparability with multiple imputation, we restrict stochasticity to Monte Carlo integration over $Z$. We also combine the stochastic approximations of $\eta$ and $Z$, using the fact that the integration is over the product density $(\eta,Z)\sim p(\eta\,|\,D_{1})\,p(Z)$. Hence, for each $\eta_{i}$ we can sample $Z_{i}\overset{\text{iid}}{\sim}p(Z)$ and $\{(\eta_{i},Z_{i})\}$ is a sample from $p(\eta\,|\,D_{1})\,p(Z)$. We then use the objective
$\hat{\mathcal{L}}_{Z}(\vartheta)=\frac{1}{N}\sum_{i=1}^{N}\big[\log|\det J_{z}T_{\vartheta}(\eta_{i},Z_{i})|+\log p\!\left(D_{2}\,|\,T_{\vartheta}(\eta_{i},Z_{i}),\eta_{i}\right)+\log p\!\left(T_{\vartheta}(\eta_{i},Z_{i})\,|\,\eta_{i}\right)\big].$ (14)

### v2

@@ 2.2 Variational Inference
Variational inference (VI) provides a practical alternative by approximating posterior distributions via optimization (Blei et al.,, 2017). Given data $D$ with model specified by some parameters $\phi$, VI seeks a distribution $q(\phi)$ within a variational family $\mathcal{Q}$ that minimizes the KLD to the posterior distribution $p(\phi\,|\,D)$, to get the optimal solution
${q}^{*}(\phi)=\underset{q(\phi)\in\mathcal{Q}}{\arg\min}\,\,\mathrm{KL}\,\!\left({q}(\phi)\parallel p(\phi\,|\,D)\right).$ (3)
If no restriction is placed on the family $\mathcal{Q}$ of distributions, then $q^{*}(\phi)$ is exactly the true posterior distribution $p(\phi\,|\,D)$. In practice, one places some parametric assumptions on the class $\mathcal{Q}$ to obtain an optimizer that approximates the true posterior distribution.
A fundamental contribution in cut-Bayes or cutting feedback methods that enabled the use of variational inference for this problem is the result (Lemma 4.1) of Yu et al., (2023). Considering the setup of Section 2.1 with $\phi=(\theta,\eta)$ and $D=(D_{1},D_{2})$, the result proved that when $\mathcal{Q}$ is restricted to all distributions $q(\theta,\eta)$ such that $q(\eta)=p(\eta\,|\,D_{1})$, the optimization in (3) yields the joint cut-posterior (S2) as the solution. Formally,
@@
Thus the cut-posterior has the nice interpretable characterization as the best approximation of the Bayes posterior (in KLD) under the constraint of no feedback from $D_{2}$ to $\eta$. In their implementation, Yu et al., (2023) used two variational approximations, decomposing $q(\theta,\eta)=q(\theta\,|\,\eta)\,q(\eta)$, using parametric families for both, i.e., $q(\theta\,|\,\eta)=q_{\lambda_{2}}(\theta\,|\,\eta)$ and $q(\eta)=q_{\lambda_{1}}(\eta)$. To estimate $q^{*}(\eta)$, the KLD to $p(\eta\,|\,D_{1})$ was minimized with respect to $\lambda_{1}$ to obtain $\hat{q}^{*}(\eta)=q_{\hat{\lambda}_{1}}(\eta)$. Then $q_{\hat{\lambda}_{1}}(\eta)$ is plugged into the second optimization (4) to estimate the minimizing parameters $\hat{\lambda}_{2}$. The cut-posterior estimate is $q_{\hat{\lambda}_{2}}(\theta\,|\,\eta)\,q_{\hat{\lambda}_{1}}(\eta)$. Similar ideas were used in Smith et al., (2025) in the context of copula models.
@@ 2.3 Normalizing Flows
Normalizing flows provide a flexible variational family by constructing complex densities through a sequence of invertible and differentiable transformations of a simple base distribution, and have been widely explored in variational inference (Rezende and Mohamed,, 2015). The basic idea of normalizing flows is as follows. If $F$ is the cumulative distribution function (cdf) of a scalar $\theta$, then there always exists a function $T$ such that $\theta\overset{d}{=}T(Z)$, where $Z$ is some base distribution. The function $T$ is explicitly obtained from the inverse cdf transformation, e.g., if $Z\sim N(0,1)$, $T(Z)=F^{-1}(\Phi(Z))$ where $\Phi$ is the standard Normal cdf. Hence, the task of approximating the distribution of $\theta$ becomes equivalent to approximating the function $T$. As neural networks are universal approximators of many function classes, it thus suffices to model $T$ using a suitable class of neural networks $g_{\lambda}$ and estimate the neural parameters (weights and biases) $\lambda$ using the variational (KL) loss. Normalizing flows build on this core idea, but the specific choices to model $T$ vary widely. We will discuss some common flow classes when introducing our method and studying its theoretical properties. A more comprehensive review of normalizing flows and their extensions can be found in (Kobyzev et al.,, 2020; Papamakarios et al.,, 2021).
@@ 3 Neural Variational Inference for Cut-Bayes
We propose a variational inference method for cutting feedback that $(a)$ can directly use samples of $\eta$ from the upstream posterior $p(\eta\,|\,D_{1})$ without requiring access to the upstream data $D_{1}$ and upstream model (likelihood $p(D_{1}\,|\,\eta)$ and prior $p(\eta)$); and $(b)$ uses normalizing flows to non-parametrically model the conditional cut-posterior of $\theta$ given $\eta$.
Recall from (4) that the cut-posterior is the solution to the variational optimization with the variational family $q(\theta,\eta)$ constrained to have $q(\eta)=p(\eta\,|\,D_{1})$ to prevent feedback. As $q(\eta)$ is fixed, we make two observations. First, the optimization in (4) is essentially only over the class of conditional distributions $q(\theta\,|\,\eta)$. Also, under this constraint of $q(\eta)=p(\eta\,|\,D_{1})$, the KL loss function in (4) can be written as
$\mathrm{KL}\,\!({q}(\theta,\eta)\parallel p(\theta,\eta\,|\,D_{1},D_{2}))=\mathbb{E}_{\eta\sim p(\eta\,|\,D_{1})}\big[\mathrm{KL}\,\!(q(\theta\,|\,\eta)\parallel p(\theta\,|\,\eta,D_{2}))\big].$
@@
Thus, finding the joint cut-posterior essentially reduces to estimating the collection of conditional distributions $\{q(\theta\,|\,\eta)\}_{\eta}$. No approximations to $p(\eta\,|\,D_{1})$ is needed. In practice, we will typically have samples $\eta_{1},\ldots,\eta_{N}$ from $p(\eta\,|\,D_{1})$. As long as the samples are iid or from an ergodic MCMC, and we have enough samples, we can approximate the expectation term in (5) with the sample average, leading to the estimate
@@
Equations (5) and (6) illustrate how the joint cut-posterior can be estimated without access to the upstream data $D_{1}$ and the model used in upstream analysis, as long as the samples $\{\eta_{i}\}$ are available. Even if $D_{1}$ and its model are available, there is no need to approximate $p(\eta\,|\,D_{1})$ using the additional variational inference step, as done in Yu et al., (2023) or Smith et al., (2025). One can run a standard MCMC, to estimate $p(\eta\,|\,D_{1})$ and use the posterior samples $\{\eta_{i}\}$ in (6). Once the downstream analysis model likelihood $p(D_{2}\,|\,\theta,\eta)$ and the prior $p(\theta\,|\,\eta)$ are chosen, the only unknown in (6) is the class of conditional distributions $q(\theta\,|\,\eta)$. In the next section, we propose using normalizing flows that can approximate any conditional family, leading to a provably accurate estimate of the joint cut-posterior.
@@ 3.1 Normalizing Flows for Conditional Distribution
We model $q(\theta\mid\eta)$ using conditional normalizing flows, which represent complex distributions as transformations of simple base distributions. Specifically, for $\theta\in\mathbb{R}^{d}$, we define a family of conditional distributions $q_{\vartheta}(\theta\mid\eta)$ corresponding to the law $\theta\mid\eta\sim T_{\vartheta}(\eta,Z)$, where $T_{\vartheta}(\eta,z)$ is a neural network-based transformation with arguments $(\eta,z)$ and parameters $\vartheta\in\Theta$, and $Z\in\mathbb{R}^{d}$ is a latent variable drawn from an easy-to-sample base distribution $p(Z)$. A common choice for $p(Z)$ is the standard normal, $Z\sim N(0,I_{d})$. However, when the true posterior exhibits heavy tails, alternative base distributions such as the Cauchy or Student’s $t$ may be more appropriate. Theoretical considerations guiding this choice are discussed in Section 4.
@@
By drawing $Z_{i}\overset{\text{iid}}{\sim}p(Z)$, we can have samples $\{\eta_{i},T_{\hat{\vartheta}}(\eta_{i},Z_{i})\}$ from this cut-posterior. We call our method NeVI-Cut (neural variational inference for cut-Bayes). Note that NeVI-Cut is completely agnostic to what upstream data $D_{1}$ and model were used to obtain the samples $\{\eta_{i}\}$ as long as averages using this set of samples approximate population averages with respect to $p(\eta\,|\,D_{1})$. It also works with any choice of base distribution, any dimension and support set of the parameters $\theta$ and $\eta$, and any class of normalizing flows. Section S2 presents some important classes of normalizing flows that we use in practice. The theoretical justification for their use is presented in Section 4.
@@ 3.2 Triply-Stochastic Algorithm
The computational challenges in calculating the objective in (8) can be three-pronged. First, a large number of samples $\eta_{i}$ may be available from the upstream analysis. Computing the NeVI-Cut ELBO, averaged over all the upstream samples, for every iteration can be slow. Second, the expectation with respect to $Z$ is typically not available in closed form. Third, the likelihood $p(D_{2}\,|\,\theta,\eta)$ might also be slow to compute especially for large datasets. We address these challenges by leveraging three sources of stochasticity.
Mini-batching or stochastic gradient descent is widely used to speed up neural network optimization. The main idea is that terms in the objective function involving averages or sums over many terms can be replaced by averages (or appropriately scaled sums) over smaller subsets (mini-batches) of these terms. Only one mini-batch is used at a time, and the mini-batches are cycled throughout the optimization. Stochastic gradient descent is an extreme case with a mini-batch size of one. Each of these three challenges has a scalable solution using such stochastic approximations. At each iteration, we can replace the average over all $N$ samples of $\eta$ with a small minibatch $\mathcal{B}_{\eta}$ of size $N_{\eta}$. Similarly, the expectation with respect to $Z$ is replaced by an average of $N_{Z}$ iid observations $\{Z_{j}\}$ sampled from $p(Z)$. The doubly stochastic approximation of (8) is thus given by
$\hat{\mathcal{L}}_{DS}(\vartheta)=\frac{1}{N_{\eta}N_{Z}}\sum_{i\in\mathcal{B}_{\eta}}\sum_{j=1}^{N_{Z}}\Big[$ $\log|\det J_{z}T_{\vartheta}(\eta_{i},Z_{j})|+$ (10)
@@
Finally, if the downstream data $D_{2}$ consists of iid units $Y_{1},\ldots,Y_{n}$, then we have $p(D_{2}\,|\,\theta,\eta)=\prod_{i=1}^{n}p(Y_{i}\,|\,\theta,\eta)$. And $\log p(D_{2}\,|\,\theta,\eta)$ can be approximated using a minibatch $\mathcal{B}_{D}$ of size $N_{D}$ as $\frac{n}{N_{D}}\sum_{k\in\mathcal{B}_{D}}\log p(Y_{k}\,|\,\theta,\eta)$. This gives the triply stochastic approximation
$\hat{\mathcal{L}}_{TS}(\vartheta)=\frac{1}{N_{\eta}N_{Z}}\sum_{i\in\mathcal{B}_{\eta}}\sum_{j=1}^{N_{Z}}\Big[$ $\log|\det J_{z}T_{\vartheta}(\eta_{i},Z_{j})|+$ (11)
$\frac{n}{N_{D}}\sum_{k\in\mathcal{B}_{D}}\log p\big(Y_{k}\,|\,T_{\vartheta}(\eta_{i},Z_{j}),\eta_{i}\big)+\log p\!\left(T_{\vartheta}(\eta_{i},Z_{j})\,|\,\eta_{i}\right)\Big].$
@@
Input: Upstream posterior samples $\{\eta_{i}\}_{i=1}^{N}$, downstream data $D_{2}$, functional form of downstream likelihood $p(D_{2}\,|\,\theta,\eta)$ and downstream prior $p(\theta\,|\,\eta)$, base distribution $p(Z)$, maximum number of iterations $S$, patience $S_{\text{patience}}$ (number of steps before early stopping), normalizing-flow-based transformation $T_{\vartheta}$ with parameter $\vartheta$ (e.g., Neural Spline Flows or Unconstrained Monotonic Neural Network Flows), learning rate $\alpha$.
@@
  Estimate ELBO:
$\hat{\mathcal{L}}_{Z}(\vartheta)=\frac{1}{N}\sum_{i=1}^{N}\big[\log|\det J_{z}T_{\vartheta}(\eta_{i},Z_{i})|+\log p\!\left(D_{2}\mid\theta_{i},\eta_{i}\right)+\log p\!\left(\theta_{i}\mid\eta_{i}\right)\big].$
@@
Output: Trained conditional flow $\hat{q}(\theta\,|\,\eta)\overset{\text{d}}{=}T_{\hat{\vartheta}}(\eta,Z),Z\sim p(Z)$, cut posterior estimate $\theta\,|\,\eta\sim\hat{q}(\theta\,|\,\eta)]\times[\eta\sim\text{Unif}\{\eta_{1},\ldots,\eta_{N}\}]$, samples from the cut-posterior $\{\eta_{i},\widetilde{\theta}_{i}\}_{i=1}^{N}$.
In practice, all three approximations are not always required simultaneously, and the level of stochasticity, i.e., choices of the minibatch sizes $N_{\eta},N_{Z},N_{D}$ can be adjusted depending on computational trade-offs. For example, if the data are not iid or not large, mini-batching of the likelihood is not recommended, and the third source can be omitted. In our experiments, to ensure comparability with multiple imputation which uses all the upstream samples $\{\eta_{i}\}$, we restrict stochasticity to Monte Carlo integration over $Z$. We also combine the stochastic approximations of $\eta$ and $Z$, using the fact that the integration is over the product density $(\eta,Z)\sim p(\eta\,|\,D_{1})\,p(Z)$. For each $\eta_{i}$ we can sample $Z_{i}\overset{\text{iid}}{\sim}p(Z)$ and $\{(\eta_{i},Z_{i})\}$ is a sample from $p(\eta\,|\,D_{1})\,p(Z)$, and use the objective
$\hat{\mathcal{L}}_{Z}(\vartheta)=\frac{1}{N}\sum_{i=1}^{N}\big[\log|\det J_{z}T_{\vartheta}(\eta_{i},Z_{i})|+\log p\!\left(D_{2}\,|\,T_{\vartheta}(\eta_{i},Z_{i}),\eta_{i}\right)+\log p\!\left(T_{\vartheta}(\eta_{i},Z_{i})\,|\,\eta_{i}\right)\big].$